OpenStudy (anonymous):

need some help here sin 7x/2. cos 3x/2 + sinx/2.cos5x/2 + sin2x.cos7x = 0

OpenStudy (psymon):

The /2 is a part of the angle, correct? Or isit sin(7x)/2 * cos(3x)/2, etc?

OpenStudy (anonymous):

it's sin(7x/2) . cos ( 3x/2) ...

OpenStudy (dumbcow):

use these identities: http://en.wikipedia.org/wiki/Trigonometric_identity#Product-to-sum_and_sum-to-product_identities it will reduce to: \[\frac{\sin (5x) +\sin(2x)}{2} + \frac{\sin (3x) -\sin(2x)}{2} + \frac{\sin (9x) -\sin(5x)}{2} = 0\]

OpenStudy (anonymous):

so i got uhm...x=kpi/6 and x=pi/6 + kpi/3 :<

OpenStudy (anonymous):

anyway thank you